OpenStudy (selrachcw95):

ok 1 min to get my question up

OpenStudy (anonymous):

you have 30 seconds left...

OpenStudy (selrachcw95):

Oh don't need to be exact.Lol!!!

OpenStudy (selrachcw95):

There u go

OpenStudy (anonymous):

ooookay. so it crosses the y axis when x = 0. when is x = 0?

OpenStudy (selrachcw95):

When it's in the middle

OpenStudy (anonymous):

what the hell am I typing? Ive gone mad. when x = 0 y = -3

OpenStudy (anonymous):

do you see?

OpenStudy (selrachcw95):

no

OpenStudy (selrachcw95):

lol

OpenStudy (anonymous):

what do we have y = (3/4)x - 3 or something

OpenStudy (selrachcw95):

srry terrible at this

OpenStudy (anonymous):

this is going to be hilarious, lets pretend that x = 0, so we now have y = (3/4)(0) - 3 which is y = -3, because (3/4) times 0 = 0

OpenStudy (anonymous):

what joy!

OpenStudy (selrachcw95):

who's joy?

OpenStudy (anonymous):

you me, we all are. because we have got y intercept!

OpenStudy (selrachcw95):

Ok I'm not getting u nehow.Lol!!!

OpenStudy (selrachcw95):

So is -3 and 0 my ansur?

OpenStudy (anonymous):

(0, -3) is your first coordinate. Your y intercept

OpenStudy (selrachcw95):

OK now what?

OpenStudy (selrachcw95):

Can u make it ne easier?

OpenStudy (anonymous):

and then to find another, that's easy peasy. Just put in a random value for x. say x = 4, nonono x = 3, i like it when x = 3. if x = 3 the equation is y = (4/3)(3) -3 y = 12/3 - 3 y = 4 - 3 y = 1 so we have (3,1) is another point!

OpenStudy (anonymous):

cant make it easier than the math baby

OpenStudy (anonymous):

just slam your head into your textbook. I promise that will work

OpenStudy (selrachcw95):

Lol!!!

OpenStudy (selrachcw95):

Ok ty for trying.

OpenStudy (anonymous):

awhhh. I failed.

OpenStudy (selrachcw95):

yea. :(

OpenStudy (selrachcw95):

It's ok.

OpenStudy (selrachcw95):

Just need a teacher.Lol!!!

OpenStudy (anonymous):

Or maybe, Just run away? You could always run away. Running is good

OpenStudy (selrachcw95):

Nah.